OpenStudy (anonymous):

The lab table has a perimeter of 214 inches. The length is 13 inches less than 3 times the width. How many inches long is the table?

OpenStudy (anonymous):

@DavidUsa

OpenStudy (anonymous):

Width: 30 Length: 77 30*3=90 90-13=77 77+77+30+30=214

OpenStudy (anonymous):

@cancherolugano